# Semantic Web technologies in sensor-based personal health monitoring systems: A systematic mapping study

### v2

@@ 18 Hussain and Park [93] 2021 Stroke Ontology ML; Queries; Rules Modular
ABP - Abnormal blood pressure; ASP - Answer set programming; BN - Bayesian network; CBR - Case-based reasoning; CDL - Contextual defeasible logic; FL - Fuzzy logic; ML - machine learning; NLP - Natural language processing.
@@ 18 Hussain and Park [93] Body (ECG); Other (health/medical records)
[…]
@@
24 Lopes de Souza et al. [99] Body (HR, BP, BT, acceleration, orientation/angular velocity); ; Other (user-submitted data)
@@
43 Zhou et al. [71] Body (BP, HR, RR, BT, SpO2, BG, uric acid, cholestrol, lipoproteins, triglycerides, sleep); Ambient (inhalable particulate matter, CO2, temperature, formaldehyde, total volatile organic compounds); Other (health/medical records)
@@
BG - blood glucose; BP - blood pressure; BT - body temperature; CO2 - carbon dioxide; CO - carbon monoxide; ECG - electrocardiogram; EEG - electroencephalogram; EMG - electromyogram; GPS - global positioning system; HR - heart rate; O2 - oxygen; RR - respiratory rate; SpO2 - blood oxygen saturation
5 Challenges in health monitoring systems
5.1 Interoperability
Interoperability can be defined as the ability of different components or systems not only to exchange information but also to make use of it [113]. There are three types of interoperability identified in the health domain: technical, semantic, and process interoperability [114, 113]. Technical interoperability refers to the way data or information moves from one system or component to another. Related to this is syntactic interoperability, which provides a structure and syntax for the transmitted data [115]. Semantic interoperability refers to the ability of the recipient to understand and make use of the received data, whereas process interoperability concerns the way in which different systems are used in actual work settings. A subset of this is clinical interoperability, through which patients can be seamlessly transferred between different care teams [113].
5.1.1 Technical interoperability
Differing data transmission technologies can contribute to a lack of technical interoperability in health monitoring systems, particularly those that use a range of different sensors. Data transmission protocols used in sensors include Bluetooth, Bluetooth Low Energy, ANT+, and Zigbee, with the first three being the most common among wearable devices today [11]. Interoperability among these different protocols can be achieved using gateway devices, which receive data from different sensors and transmit it to cloud services [116]. This is done by Ali et al. [79], who use a router as a gateway to receive sensor data and transmit it to the internet. A number of the systems [78, 80, 84, 85, 93, 96, 99, 104, 109, 112] use a mobile phone as a gateway device or base station, typically receiving sensor data via Bluetooth or Bluetooth Low Energy and transmitting it to the cloud via Wi-Fi or mobile data.
5.1.2 Syntactic interoperability
While technical interoperability is associated with hardware components and infrastructure, syntactic interoperability is usually associated with data formats [117]. There are several standards that are widely used to promote syntactic interoperability among systems. Among them is the ISO/IEEE 11073 standard, which provides a common format for communication involving medical devices and patient health data, with an emphasis on vital signs. This is used by El-Sappagh et al. [84] for message formatting between body sensors and the base unit. Other important standards for health data are provided by Health Level 7 (HL7). One of these is Fast Health Interoperability Resources (FHIR), which describes data formats, resources, and an application programming interface (API) through which health information can be exchanged [113]. El-Sappagh et al. [84] integrate FHIR in their proposed system, converting sensor data from the ISO/IEEE 11073 standard to FHIR resource formats. Additionally, the system receives data in FHIR format from hospital information systems. In this way, both sensor data and data from hospital systems are in the same format. FHIR resources can be defined using different data formats [^21] https://build.fhir.org/resource-formats.html, including XML, JavaScript Object Notation (JSON), and Terse RDF Triple Language (Turtle).
5.1.3 Semantic interoperability
The next type of interoperability is semantic interoperability, which is concerned with the meaning of the exchanged information. Semantic interoperability can be achieved through the use of unambiguous codes and identifiers, which can be provided by existing standard classifications and terminologies [113]. Ontologies are, of course, a well-established way to embed semantic interoperability in a system [118]. Within the medical domain, many existing medical terminologies are available as ontologies, including SNOMED CT [^22] https://bioportal.bioontology.org/ontologies/SNOMEDCT, the International Classification of Diseases (ICD) [^23] https://bioportal.bioontology.org/ontologies/ICDO, and the International Classification for Nursing Practice (ICNP) [^24] https://bioportal.bioontology.org/ontologies/ICNP. Among the systems, SNOMED CT is the most commonly used [81, 84, 98, 99, 73, 105, 107, 71]. ICNP is used by Elhadj et al. [85] and Henaien et al. [90] , while ICD is used by Spoladore et al. [106] and Yu [70] (ICD-11, the latest version) as well as Titi et al. [107] (ICD-10). The Unified Medical Language System (UMLS) [119] is a large thesaurus that integrates multiple terminologies of medical knowledge. It is used by Peral et al. [104] and Zhou et al. [71]. Another thesaurus is Medical Subject Headings (MeSH), which is used for indexing, cataloguing, and searching health information, and is integrated in the system proposed by Reda et al. [73]. Spoladore et al. [106] incorporate the International Classification of Functioning, Disability and Health (ICF) [^25] https://icd.who.int/dev11/l-icf/en.
[…]
Semantic Web technologies also provide a means to represent sensors and the data they capture. Sensors can be represented with varying degrees of expressiveness. Concepts that can be captured about sensors include unique identifier, manufacturer, location of deployment, dimensions, operating conditions, type of data captured, and hierarchy with regard to related sensors [59]. Similarly, various sensor data concepts can be represented, such as the property being observed, units of measurement, and measurement timestamps. Most systems represent sensor and sensor data concepts in ontologies, with 13 reusing and extending existing sensor or device ontologies, namely SSN/SOSA [81, 84, 85, 94, 105, 74, 107], SAREF and its extensions [83, 89, 99, 75], the Amigo device ontology [92], and the Moving Objects ontology [105]. The reuse of existing sensor ontologies, particularly established ones such as SAREF, can contribute to a higher degree of expressiveness for sensor and sensor data concepts. This is because these validated ontologies provide rich modelling of such concepts, facilitating more effective querying of and reasoning on sensor data, which is essential for situation analysis. Comprehensive sensor ontologies also support sensor management, allowing sensors to be catalogued based on their attributes as captured in ontologies [59].
Foundational ontologies can contribute to semantic interoperability by providing unambiguous and domain-independent concept definitions [120]. Three of the selected systems directly incorporate a foundational ontology. El-Sappagh et al. [84] use the Basic Formal Ontology, while De Brouwer et al. [83] and Stravropoulos et al. [74] use the DOLCE+DnS (Description and Situation) Ultra Lite (DUL) ontology. Other systems indirectly integrate foundational ontologies by reusing other ontologies that have already incorporated them. For example, the SSN ontology uses DUL as its upper ontology [19], and the SAREF ontology also has an indirect reference to DUL through its mappings to the SSN ontology [121]. Consequently, any system that reuses the SSN or SAREF ontologies inherits an indirect connection to DUL.
5.1.4 Process interoperability
The final type of interoperability is process interoperability, which focuses on how systems and components work seamlessly together in real-world settings. One way to enhance process interoperability in health monitoring systems is through the integration of sensor data with comprehensive health and medical records [114]. The inclusion of these records allows health monitoring systems to complement and extend healthcare provided in clinical settings. Health and medical records provide additional information that is useful for health monitoring, such as an individual’s disease history, laboratory test results, medications taken, allergies, and previous hospital admissions. 19 systems integrate existing records in some way, with most of them represented using ontologies. The systems proposed by Ali et al. [78, 79], El-Sappagh [84], and Rhayem et al. [105] have the most comprehensive records, capturing laboratory tests, prior disease diagnoses, and lifestyle information such as exercise, nutrition, alcohol consumption, and smoking status. Some systems use medical records to extract diagnosis status [77], while others use them to extract an individual’s risk factors for disease [78]. These records can also be used to overcome limitations of sensor data such as missing values, as was done by Ali et al. [77]. Besides health and medical records, data from social networks and other web and mobile applications can also be used to complement sensor data. For instance, Ali et al. [77] use social networking data for monitor individuals’ mental health through sentiment analysis.

### v3

@@ 21 Hussain and Park [120] 2021 Stroke User validated Ontology ML; queries; rules Modular
ASP - Answer set programming; BN - Bayesian network; CBR - Case-based reasoning; CDL - Contextual defeasible logic; COPD - Chronic obstructive pulmonary disease; FL - Fuzzy logic; KG - knowledge graph; ML - machine learning; NLP - Natural language processing
5 Challenges in health monitoring systems
5.1 Interoperability
Interoperability can be defined as the ability of different components or systems not only to exchange information but also to make use of it [134]. There are three types of interoperability identified in the health domain: technical, semantic, and process interoperability [135, 134]. Technical interoperability refers to the way data or information moves from one system or component to another. Related to this is syntactic interoperability, which provides a structure and syntax for the transmitted data [136]. Semantic interoperability refers to the ability of the recipient to understand and make use of the received data [134], whereas process interoperability concerns the seamless coordination of workflows in care delivery [137]. A subset of this is clinical interoperability, through which patients can be seamlessly transferred between different care teams [134]. This review focuses on technical, syntactic, and semantic interoperability. We also discuss the role of interoperability in enabling data fusion, a crucial functionality in health monitoring systems.
5.1.1 Technical interoperability
Differing data transmission technologies can contribute to a lack of technical interoperability in health monitoring systems, particularly those that use a range of different sensors. Data transmission protocols used in sensors include Bluetooth, Bluetooth Low Energy, ANT+, and Zigbee, with the first three being the most common among wearable devices today [18]. Interoperability among these different protocols can be achieved using gateway devices, which receive data from different sensors and transmit it to cloud services [138]. This is done by Ali et al. [112], who use a router as a gateway to receive sensor data and transmit it to the internet. 11 of the systems [108][111, 113, 116, 89, 120, 121, 124, 130, 102, 133] use a mobile phone as a gateway device or base station, typically receiving sensor data via Bluetooth or Bluetooth Low Energy and transmitting it to the cloud via WiFi or mobile data.
5.1.2 Syntactic interoperability
While technical interoperability is associated with hardware components and infrastructure, syntactic interoperability is usually associated with data formats [139]. There are several standards that are widely used to promote syntactic interoperability among systems. Among them is the ISO/IEEE 11073 standard, which provides a common format for communication involving medical devices and patient health data, with an emphasis on vital signs. This is used by El-Sappagh et al. [116] for formatting data for transmission to the base unit or gateway. Other important standards for health data are provided by Health Level 7 (HL7). One of these is FHIR, which describes data formats, resources, and an application programming interface (API) through which health information can be exchanged [134]. Two of the 48 systems make use of FHIR. El-Sappagh et al. [116] convert sensor data from the ISO/IEEE 11073 standard to FHIR resource formats, while also receiving data in FHIR format from hospital information systems. In this way, both sensor data and data from hospital systems are in the same format. Similarly, Kilintzis et al. [97] propose a semantic model based on FHIR, using its data types and defining classes as FHIR resource categories. FHIR resources can be defined using different data formats [^24] https://build.fhir.org/resource-formats.html, including XML, RDF serialised in Turtle, and JSON. Both systems use the JSON format.
5.1.3 Semantic interoperability
The next type of interoperability is semantic interoperability, which is concerned with the meaning of the exchanged information. Semantic interoperability can be achieved through the use of unambiguous codes and identifiers, which can be provided by existing standard classifications and terminologies [134]. Ontologies are, of course, a well-established way to embed semantic interoperability in a system [140]. Within the medical domain, many existing medical terminologies are available as ontologies, including SNOMED CT [^25] https://bioportal.bioontology.org/ontologies/SNOMEDCT, the International Classification of Diseases (ICD) [^26] https://bioportal.bioontology.org/ontologies/ICDO, and the International Classification for Nursing Practice (ICNP) [^27] https://bioportal.bioontology.org/ontologies/ICNP. Among the systems, SNOMED CT is the most commonly used [87, 97] [114, 116, 123, 124, 106, 83, 100, 103]. ICNP is used by Elhadj et al. [89] and Henaien et al. [93] , while ICD is used by Spoladore et al. [98] and Yu [104] (ICD-11, the latest version) as well as Titi et al. [100] (ICD-10). The Unified Medical Language System (UMLS) [141] is a large thesaurus that integrates multiple terminologies of medical knowledge. It is used by Peral et al. [130], and Zhou et al. [103]. Another thesaurus is Medical Subject Headings (MeSH), which is used for indexing, cataloguing, and searching health information, and is integrated in the system proposed by Reda et al. [106]. Garcia Moreno et al. [91] and Spoladore et al. [98] incorporate the International Classification of Functioning, Disability and Health (ICF) [^28] https://icd.who.int/dev11/l-icf/en.
Terminologies for specific diseases and conditions also exist. For example, Ali et al. [110] and El-Sappagh et al. [116] reuse ontologies specific to diabetes. Similarly, De Brouwer et al. [115] use the third edition of the International Classification of Headache Disorders (ICHD-3) [^29] https://ichd-3.org, while Hristoskova et al. [119] and Zafeiropoulos et al. [107] reuse the Heart Failure Ontology and Parkinson and Movement Disorder Ontology respectively. The Vital Sign Ontology is extended by El-Sappagh [116] and Ivașcu and Negru [95]. Xu et al. [105] posit that it is difficult to build scalable ontology-based systems suitable for large amounts of healthcare data and instead opt for a linked data approach to add semantic information to the data. Their proposed system uses linked open data medical knowledge graphs, namely Diseasome, DBpedia, and DrugBank. Using these resources, they create a knowledge graph showing the relationships between symptoms and diseases. Domain-independent concepts can also be referenced from Semantic Web technologies. For instance, Peral et al. [130] and Reda et al. [106] both use WordNet, a lexical English language database of semantic relations between words, linking them into semantic relations.
Semantic Web technologies also provide a means to represent sensors and the data they capture. Sensors can be represented with varying degrees of expressiveness. Concepts that can be captured about sensors include unique identifier, manufacturer, location of deployment, dimensions, operating conditions, type of data captured, and hierarchy with regard to related sensors [74]. Similarly, various sensor data concepts can be represented, such as the property being observed, units of measurement, and measurement timestamps. 38 systems represent sensor and sensor data concepts in ontologies. The reuse of existing sensor ontologies, particularly established ones such as SAREF, can contribute to a higher degree of expressiveness for sensor and sensor data concepts. This is because these validated ontologies provide rich modelling of such concepts, facilitating more effective querying of and reasoning on sensor data, which is essential for situation analysis. Comprehensive sensor ontologies also support sensor management, allowing sensors to be catalogued based on their attributes as captured in ontologies [74]. Despite these benefits, only 16 systems reuse existing sensor or device ontologies, namely SSN/SOSA [87, 91, 125] [114, 116, 89, 95, 83, 99, 100], SAREF and its extensions [115, 118, 124, 107], the Amigo device ontology [119], and the Moving Objects ontology [83]. This could be attributed to the existing ontologies providing more complexity than the systems require, although this can be mitigated by selectively importing only the relevant classes.
Foundational ontologies can contribute to semantic interoperability by providing unambiguous and domain-independent concept definitions [142]. Three of the selected systems directly incorporate a foundational ontology. El-Sappagh et al. [116] use the Basic Formal Ontology, while De Brouwer et al. [115] and Stavropoulos et al. [99] use the DOLCE+DnS (Description and Situation) Ultra Lite (DUL) ontology. Other systems indirectly integrate foundational ontologies by reusing other ontologies that have already incorporated them. For example, the SSN ontology uses DUL as its upper ontology [29], and the SAREF ontology also has an indirect reference to DUL through its mappings to the SSN ontology [143]. Consequently, any system that reuses the SSN or SAREF ontologies inherits an indirect connection to DUL.
5.1.4 Supporting data fusion
For real-time health monitoring, streaming sensor data must be retrieved and dynamically fused with other heterogeneous, multimodal, and distributed sources of data. This data fusion is pivotal for downstream situation detection, situation prediction, and decision support. Interoperability serves as an enabler for effective data fusion. Technical interoperability provides the protocols and hardware necessary to collect data from diverse sources; syntactic interoperability ensures data formats and structures are compatible across systems; and semantic interoperability establishes shared meaning for data elements, allowing accurate interpretation. Additionally, data fusion supports process interoperability by creating a unified and comprehensive patient view from diverse data sources, which enables different healthcare providers to coordinate their workflows effectively [137].
As seen in Table LABEL:supported-data, the selected systems support a wide range of heterogeneous and multimodal data. All 48 systems collect physiological data from body sensors, while 15 systems additionally incorporate weather data from ambient sensors. Health and medical records are the most frequently incorporated data source other than sensor data, with 21 systems supporting the incorporation of such data from external hospital systems. These records provide additional information that is useful for health monitoring, such as an individual’s disease history, laboratory test results, medications taken, allergies, and previous hospital admissions. The systems proposed by Ali et al. [111, 112], El-Sappagh [116], and Rhayem et al. [83] have the most comprehensive records, capturing laboratory tests, prior disease diagnoses, and lifestyle information such as exercise, nutrition, alcohol consumption, and smoking status. Some systems use medical records to extract diagnosis status [110], while others use them to extract an individual’s risk factors for disease [111]. These records can also be used to overcome limitations of sensor data such as missing values, as was done by Ali et al. [110]. Besides health and medical records, data from social networks can also be used to complement sensor data. This is done in two systems: Ali et al. [110], who use social networking data for monitor individuals’ mental health through sentiment analysis; and Ammar et al. [108], who use social networks, blogs, and news articles as sources of public knowledge.
A detailed analysis of the data sources, including sensor devices and existing datasets, is provided in Section 6.1.
@@ 5.1.4 Supporting data fusion
6 Ammar et al. [108] Body (BP, activity); Other (health/medical records, social determinants of health, social networks, blogs)
7 Bampi et al. [87] Body (HR, BR); Ambient (Humidity, other unnamed)
@@
15 Garcia-Moreno et al. [91] Body (BT, HR, accelerometer, electrodermal activity); Ambient (unnamed)
@@ 21 Hussain and Park [120] Body (ECG); Other (health/medical records)
22 Ivașcu and Negru [95] Body (HR, RR, ECG, acceleration)
23 Ivașcu et al. [96] Body (EEG, acceleration); Ambient (video, audio, motion, bed sensor data)
@@ 22 Ivașcu and Negru [95] Body (HR, RR, ECG, acceleration)
25 Kilintzis et al. [97] Body (BP, ECG, EIT, SpO2); Other (health/medical records)
@@
28 Lopes de Souza et al. [124] Body (HR, BP, BT, acceleration, orientation/angular velocity); Other (user-submitted data)
29 Martella et al. [125] Body (HR, temperature); Ambient (CO2, TVOC, particulate matter, temperature, humidity, ambient light, atmospheric pressure, sound pressure, electromagnetic emission); Other (GPS)
@@
48 Zhou et al. [103] Body (BP, HR, RR, BT, SpO2, BG, uric acid, cholestrol, lipoproteins, triglycerides, sleep); Ambient (particulate matter, CO2, temperature, formaldehyde, TVOC); Other (health/medical records)
@@
BG - blood glucose; BP - blood pressure; BT - body temperature; CO2 - carbon dioxide; CO - carbon monoxide; ECG - electrocardiogram; EEG - electroencephalogram; EIT - electroimpedance tomography; EMG - electromyogram; GPS - global positioning system; HR - heart rate; O2 - oxygen; RR - respiratory rate; SpO2 - blood oxygen saturation; TVOC - total volatile organic compounds